Encyclopedia Astronautica
2008.04.20 - ISS On-Orbit Status 04/20/08


Ahead: Week 1 of Increment 17 (with CDR Sergei Volkov, FE-1 Oleg Kononenko, FE-2 Garrett Reisman).

Crew wake/sleep cycle has 'normalized', Today's wakeup - 2:00am, sleeptime - 5:30pm EDT.

First activity this morning for FE-2 Reisman was to start on his FD30 (Flight Day 30) session with the NASA/JSC experiment NUTRITION w/Repository. Reisman completed the all-day session, collecting urine samples for 24 hrs (to continue through tomorrow morning) and blood samples (for Serum & Heparin). (Garrett performed self-phlebotomy, i.e., drew his blood samples (from an arm vein) which were first allowed to coagulate in the Repository, then spun in the HRF2 RC (Human Research Facility 2/Refrigerated Centrifuge) and finally placed in MELFI (Minus-Eighty Laboratory Freezer for ISS). The RC was later powered off after a temperature reset to limit wear on the compressor, and cleaned. Background: NUTRITION is the most comprehensive in-flight study done by NASA to date of human physiologic changes during long-duration space flight; this includes measures of bone metabolism, oxidative damage, nutritional assessments, and hormonal changes. The Clinical Nutritional Assessment profile currently required on all U.S. Astronauts collects blood and urine samples preflight and postflight. NUTRITION expands this protocol by also capturing in-flight samples and an additional postflight sample. Furthermore, additional measurements are included for samples from all sessions, including additional markers of bone metabolism, vitamin status, and hormone and oxidative stressor tests. The results will be used to better understand the impact of countermeasures (exercise and pharmaceuticals) on nutritional status and nutrient requirements. The Clinical Nutritional Assessment profile (MR016L), first started on two Mir crewmembers and then on all ISS US crews, nominally consists of two pre-flight and one post-flight analysis of nutritional status, as well as an in-flight assessment of dietary intake using the FFQ (Food Frequency Questionnaire). The current NUTRITION project has expanded MR016L testing in three ways: Addition of in-flight blood & urine collection (made possible by MELFI), normative markers of nutritional assessment, and a return session plus 30-day (R+30) session to allow evaluation of post-flight nutrition and implications for rehabilitation.)

Later, the FE-2 worked in the JLP (Japanese Experiment Module Experiment Logistics Module Pressurized Section), performing the periodical status and shell temperature check from the MKAM (Minimum Keep-Alive Monitor) and calling down the temperatures via S-band.

CDR Volkov conducted the routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the Service Module (SM), including ASU toilet facilities systems/replaceables and the weekly collection of the toilet flush (SP) counter and water supply (SVO) readings for calldown to TsUP. (Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, replacement of the KOV EDV at the SKV-2 air conditioner for the Elektron-intended water, and processing U.S. condensate water as it becomes available in a filled CWC from the Lab humidifier.)

Sergei also gathered weekly data on total operating time & 'On' durations of the Russian POTOK-150MK (150 micron) air filter unit of the SM's SOGS air revitalization subsystem for reporting to TsUP.

At ~8:05am EDT, the crewmembers held their regular WPC (Weekly Planning Conference) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), RED resistive exercise device (FE-2) and VELO bike with bungee cord load trainer (CDR, FE-1).

Afterwards, Garrett downloaded the crew's exercise data file to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

All three crewmembers had their weekly PFCs (Private Family Conferences) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop), Sergei at 7:00am, Oleg at 8:30am, Garrett at 11:35am.

ATV Reboost Test: A single-burn reboost test of the ATV 'Jules Verne' is scheduled for tonight at 12:10am-1:52am. ISS attitude control authority will be handed over to the Russian MCS (Motion Control System) thrusters at ~12:00am and returns to US momentum management at ~2:47am.

Weekly Science Update (Expedition Sixteen -- Week 26)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Completed. Instrument was relocated to FGB for Inc17 measurements (as Russian experiment) and ALTCRISS was re-activated on 4/8. Increment 16 passive dosimeters and memory cards were returned on 15S.

ANITA: Completed.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): In progress.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): The CSLM-2 SPU #4 completed the vacuum vent cycles, and started the sample 10-hour heat soak, at the end of 10 hours the quench was successfully completed. The SPU #4 temperature data was reviewed and preliminary results show a nominal run.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Experiment data were returned on 15S.

EuTEF (European Technology Exposure Facility): EuTEF platform is nominal. DEBIE-2: Link error still under investigation; DOSTEL: On-going science acquisition; EuTEMP: Currently inactive as planned; EVC: So far further troubleshooting is required, as the high-resolution downlinked images are very difficult to interpret. The impact of MISSE-6 on the EVC field of view is apparently much bigger than anticipated; EXPOSE: On-going science acquisition; FIPEX: Short-duration measurements with sensors RAM1, RAM4 and ZENITH8 continue. Long-term measurements can not be started due to RS422 Link Error - Troubleshooting on-going; MEDET: On-going science acquisition, running nominal, except sample wheel of the Spectrometer is blocked on "park" position; PLEGPAY: During the SOLACES instrument operation, some countermeasure Xenon gas release was implemented (nominal procedure). Analysis on-going to assess if we can relax this countermeasure procedure. Next test (Langmuir Probe acquisitions) via direct commanding will take place during the 15S undocking; TRIBOLAB: on 4/10 the instrument was commanded in Thermal Stabilisation mode in preparation of the first experiment run. On 4/11, the first TRIBOLAB Pin On Disk (POD1) was started and ran until 15S undock. The experiment was paused during the 15S thruster firing on 4/14, and resumed successfully. On 4/15, the Pin On Disk (POD1) run was suddenly stopped after we encountered much higher than anticipated friction coefficients measurements. Analysis on-going.

Fluid Science Laboratory (FSL): The FSL Facility awaits further troubleshooting after 1J/A departure.

GEOFLOW: Deferred. Start of GEOFLOW is pending further FSL troubleshooting/commissioning activities.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Second session for Yuri Malenchenko has been nominally performed between 4/12 and 4/15. Blood and urine samples were returned on 15S.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.

NUTRITION/REPOSITORY: In progress.

PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Last crew sampling session for Inc16 FE-1 Malenchenko was successfully performed on 4/7. Samples were downloaded on 15S.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): In progress.

SOLAR (Solar Monitoring Observatory): On 4/12, the tracking mode software was successfully patched. First results are promising, but require further assessment. SOLAR platform went to stand-by mode during Sun tracking measurements. The Sun observation is now closed (Beta angles higher than 22deg), and will resume on 4/30. SOVIM: Science data acquisition has been performed nominally until closing of Sun observation window (4/15); SOLSPEC: One spectrum measurement has been successfully performed on 4/13 and will help to cross-compare the results gathered with Sun instrument on NASA Sounding Rocket launched on 4/14/16:58GMT; SOLACES: Two spectrum measurements have been successfully performed on 4/13 and will help to cross-compare the results gathered with Sun instrument on NASA Sounding Rocket launched on 4/14/16:58GMT.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: Planned.

WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Planned.

CEO (Crew Earth Observation): Through 4/15 the ground has received a total of 18,267 ISS CEO images for review and cataloging. Late this week imagery received included sessions with camera times corresponding to the following daily target requests: Pilcomayo River dynamics, N. Argentina; Lima, Peru; East Venezuelan land use; Irrawaddy River Delta, Burma; Chongqing, China; Luquillo Forest, Puerto Rico; Santa Barbara Coast, California; Eastern Tien Shan Mountains, China; and Slate Islands Impact Crater, Lake Superior. We will be providing feedback on these acquisitions as we work through them in the coming week. Feedback on last week's acquisitions includes: Teide Volcano, Canary Islands - first review indicates more clouds present then we had hoped; Volcán Colima, Mexico - again clouds prevented a good acquisition; imagery for other acquisitions are still under review. 'Your recent striking image of the Cananea Copper Mine, Sonora-Mexico will be published on NASA/GSFC's Earth Observatory site this weekend. Your photo beautifully illustrates the land use and ecological impacts one of the world's largest open-pit copper mines. Thanks for spotting this one. We are very pleased with the quality of your imagery and your high level of response to our target requests. Please don't hesitate to ask if we can provide additional products or assistance in acquiring our science targets.'

No CEO photo targets uplinked for today.

More... - Chronology...

Home - Browse - Contact
© / Conditions for Use